OpenStudy (anonymous):

Suppose that you were saving money over 5 years to use in a purchase later. You have $1000 to put in the savings. After surveying several banks for savings plans, you found these options. A stands for the amount you will have in the bank after x years.

OpenStudy (misssunshinexxoxo):

A = 1000 + 1000(0.05)5. A = 1000 + 250 A = 1250 So A is a lower value than D. B = (1000)(2.71828)^(0.05 * 5) B = 1000 * (2.781828^.25) B = 1284.03 So B is more than A or D. The sequence now is B, D, A. C = (1000)(1.05)^5. C = 1276.28 So the sequence of values is B, C, D, A.

OpenStudy (misssunshinexxoxo):

Yes :)

OpenStudy (anonymous):

ok and then Does the highest amount make the most sense to you? Explain your point of view. @misssunshinexxoxo

OpenStudy (misssunshinexxoxo):

Of course; by plugging in the calculations we verified it is correct!

OpenStudy (misssunshinexxoxo):

My pleasure :)